OpenStudy (anonymous):

how do i solve x^2+9x+20/x^2-25 divided by x+4/x-4

OpenStudy (anonymous):

\(\huge \frac{x^2+9x+20}{x^2-25}\times \frac{x-4}{x+4}=\frac{(x+4)(x+5)}{(x+5)(x-5)}\times \frac{x-4}{x+4}\\ \huge=\color{blue}{\frac{x-4}{x-5}}\)